OpenStudy (prepfortears):

Find all solutions in the interval [0, 2π). 2 sin^2(x) = sin(x)

OpenStudy (prepfortears):

I will award a medal

OpenStudy (campbell_st):

set it to zero so 2sin^2(x) - sin(x) = 0 factorise sin(x)[2sin(x) -1] = 0 you should be able to solve it from here

OpenStudy (anonymous):

Well, let's see. You want to find all the solutions for: $$ 2sin^2(x) = sin(x) $$We could simplify that by dividing by sin(x). The problem is that the division is defined only when \(sin(x) \ne 0\). So let's first check what happens when \(sin(x) = 0\): $$ 2(0)^2 = 0 \\ 0 = 0 $$ Means, when \(sin(x) = 0\) we have a solution. That means that: $$ x = \pi\cdot K $$ are some of the solutions of the equation. Now, in case \(sin(x) \ne 0\) we can divide: $$ 2sin^2(x) = sin(x) \\ \frac{2sin^{\cancel2}(x)}{\cancel{sin(x)}} = \frac{\cancel{sin(x)}}{\cancel{sin(x)}} \\ 2sin(x) = 1 \\ sin(x) = \frac{1}{2} $$Means, when \(sin(x) = \frac{1}{2}\) we also have solutions for the equation. $$ x = \frac{\pi}{6} + 2\pi \cdot K \\ x = \frac{5\pi}{6} + 2\pi \cdot K $$ Limiting the solutions to the given interval: $$ 0 \le x \lt 2\pi \\ x = 0,\; x = \pi \;\;\;\;\;\;\;\;\;\text{for } sin(x) = 0 \\ x = \frac{\pi}{6},\; x = \frac{5\pi}{6}\;\;\;\;\text{for } sin(x) \ne 0 \\ $$ And we can also verify: http://www.wolframalpha.com/input/?i=2sin%5E2%28x%29+%3D+sin%28x%29%2C+0%3C%3Dx%3C2pi